# Temporal Drift in Privacy Recall: Users Misremember From Verbatim Loss to Gist-Based Overexposure Over Time

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Social aspects of security and privacyccs: Human-centered computing Empirical studies in HCI
@@ 1.2. Problem and Stakes
People reshare or cross-post older content all the time. The audience selected at creation might be forgotten months later as context changes. This gap throughtemporal drift in privacy recall opens up unintended disclosure risks, complementing classic context collapse (danah boyd and Marwick, 2011; Vitak, 2012). We focus on a ubiquitous visual cue: perceived age-and draw from it design implications for making past audiences recognizable at the moment of reuse.
@@ 1.3. Operational Definition and Threat Model
Figure 1. Temporal drift schematic: recall specificity declines as perceived age increases, raising the chance of scope misremembering.:
@@ 1.4. Scope and Assumptions
  • RQ1: How does a system capture recall and convert it to recognition with very little friction?
  • RQ2: How do teams conduct the evaluation of such patterns without going out and collecting usability data?
  • RQ3: Why does the perceived age plausibly reduce recall of past audiences?
  • RQ4: And when people err, why do the errors tend more toward the elder audiences?
@@ 2. Related Work
In the literature on social media, the audience reasoning is entwined with the idea of an imagined audience: a poster forms models of the sort of people who might see their content and then behave accordingly-if those mental models are often miscalibrated (danah boyd and Marwick, 2011; Vitak, 2012; Marwick and boyd, 2011; Litt and Hargittai, 2016; Bernstein et al., 2013). Due to context collapse, heterogeneous groups of people-family, friends, work colleagues, etc.-are compressed into single broadcast venues, which makes it difficult to maintain types of scopes that must be occasionally nuanced. Even if technological aids are in play, e.g., friends lists, groups, restricted audiences, they demand one to keep up their attention just in time, often at the cost of gradual degradation due to changes in membership or defaults. It is, thus, a social as well as interface problem: the UI lets one choose to emphasize certain dimensions (such as elapsed time) and glaringly omit prior audience, leading to the formation of misinformed choices and decisions (Rader, 2014; Eslami et al., 2016).
Contextual integrity frames privacy as appropriate information flow—what attributes, transmitted under which principles, to which roles, in which contexts (Nissenbaum, 2004). Audience provenance—“who could see this then”—is a core parameter of that flow. When resurfacing elides provenance, users must infer appropriateness without a key dimension. That omission is consequential when content migrates across apps with incompatible audience ontologies or when “memories” features highlight time but not the original recipients (Gilbert et al., 2023; Brandtzaeg and Lüders, 2018). Empirically, small changes in defaults and salience can shift disclosure (Wang et al., 2014; Stutzman and Kramer-Duffield, 2010; Madejski et al., 2012); in reverse, omission of salient parameters can yield miscalibrated behavior.
@@
Imagined & invisible audiences People misestimate who sees posts; mixed groups collapse into one venue (danah boyd and Marwick, 2011; Litt and Hargittai, 2016; Bernstein et al., 2013). Make the actual prior audience visible at reuse time to correct mental models.
Contextual integrity Privacy = appropriate flows given roles & transmission principles (Nissenbaum, 2004). Surface “who could see this then” as first-class provenance during reuse.
Memory dynamics Verbatim fades faster than gist; source misattribution (Ebbinghaus, 1913; Johnson et al., 1993; Schacter, 1996). Convert recall to recognition; show prior scope, not just age.
Choice architecture Small cues & defaults steer disclosure (Wang et al., 2014; Johnson and Goldstein, 2003; Thaler and Sunstein, 2008). Use narrow-first defaults when provenance is missing; avoid dark patterns.
@@
Memory research makes the temporal dimension sharper. Verbatim details decay rapidly compared to gist (Ebbinghaus, 1913; Schacter, 1996), and people often commit source-monitoring errors—remembering that they shared but misattributing audience or channel (Johnson et al., 1993; Brainerd and Reyna, 2002; Roediger and McDermott, 1995; Schacter et al., 2007). When a large “6 months ago” label is visible and audience provenance is not, retrieval relies on habits (“I usually post publicly”) or norms (“family photos are for friends”), leading to one-step drift along typical audience orders. Photography and platform automation can further distort memory; for instance, offloading memory to photos or feeds changes what details people recall (Henkel, 2014; Sparrow et al., 2011). Overconfidence compounds the problem: people can feel certain while being wrong (Koriat, 2012; Moore and Healy, 2008).
There need to be three ingredients for temporal drift: the salience of age, invisibility of provenance, and an audience lattice that allows for ”one step more open” errors. Further, some remedies are suggested: Display the relevant parameters at the point of decision (Egelman et al., 2010); introduce ethical defaults (Johnson and Goldstein, 2003; Thaler and Sunstein, 2008); eschew dark patterns (Gray et al., 2018; Mathur et al., 2019); and help people form good folk theories by using outcome-based copy, not technical jargon (Rader, 2014). We take this into concrete flora of patterns and a risk-aware policy for when to show them.
@@ 3.1. Mechanism & Core Patterns
From the standpoint of temporal drift, it arises as a boundary-case within the recognition-versus-recall framework. If we think about the resurfacing of an old item as a recall event, then it is…” For years, human-computer interaction wisdom has advised favoring recognition over recall because ”recognition means less brain power and fewer chances to go wrong” (Nielsen, 1994; Sweller, 1988). That seemingly innocent timestamp devoid of provenance violates that principle just when it should have been upheld (Nissenbaum, 2004; Rader, 2014; Eslami et al., 2016). The temporary fix is simple: pass the audience parameter so that the user interface offers recognition with possible deferral (Moreau and Missier, 2013; Egelman et al., 2010). It endorses autonomy as an expression from sharing while making sure there is no unwanted overexposure.
@@
A first pattern is a portable provenance badge: a short, readable label that travels with the media wherever it appears in reuse flows, for example, “Original audience: Friends — May 2024.” The badge answers the precise, situated question posters are likely to have (“who could see this then?”) and can be implemented as metadata read at render time (Nissenbaum, 2004; Bernstein et al., 2013; danah boyd and Marwick, 2011; Moreau and Missier, 2013). Tapping reveals a small, accessible sheet that preselects “Carry over Friends” and provides one-tap override to broaden or narrow (Johnson and Goldstein, 2003; Thaler and Sunstein, 2008). The interaction is designed for recognition, not search; it keeps routine reuse fast while making risk-relevant choices reversible and explicit (Nielsen, 1994).
@@
A second pattern involves the context-restoration prompt, which surfaces only when it is needed most. Many platforms, ranging from WeChat to Instagram, can locally infer indecision; for instance, by detecting on-device hesitation or backtracking through the interface, but without logging raw behavior (Tsai et al., 2011; Leon et al., 2012). When an item is marked as Old, and uncertainty is detected - or if the content category is sensitive - the UI jumps up briefly to confirm its prior audience with a narrow-first alternative (Brandtzaeg and Lüders, 2018). The copy never scolds or talks above the user in jargon but instead describes outcomes (”only your friends will see this”) (Rader, 2014; Eslami et al., 2016). When user certainty appears to be very high, the UI stays silent unless the requested scope conflicts with provenance, in which case a gentle clarification pops up, with an easy option to preserve the original scope (Koriat, 2012; Moore and Healy, 2008).
@@
The third pattern comes in for imports without provenance. Cross-app reuse frequently strips or reinterprets audience information (Gilbert et al., 2023). In these situations, it is more ethical to start with the conservative scope while providing an easy way to broaden it than to widen the scope silently. Defaults do possess power; hence, they should be used to defend a user’s rights, not to prey on them (Johnson and Goldstein, 2003; Thaler and Sunstein, 2008). When audience mappings turn out to be lossy or ambiguous, the interface should make the ambiguity explicit and let the poster decide (Gray et al., 2018; Mathur et al., 2019).
@@ 3.2. Accessible Deployment, Portability, & Governance
Those patterns have to be accessible and international in implementation. Screen-reader labels should expose the badge text (e.g., ‘Original audience: Friends‘) and the purpose for controls (e.g., ‘carry over original audience‘). Labels, icons, and focus order must allow keyboard navigation; color contrast must consider common impairments; date formats must be localized; and right-to-left layouts need validation. Copies are outcome-oriented, not mechanism-oriented. They say what will happen (‘only your friends will see this‘) and not how the system judges it. Because users build folk theories about feeds and privacy controls (Eslami et al., 2016; Rader, 2014), correct and easy-to-understand explanations reduce the confusion and increase the users’ trust.
The patterns are also portable beyond social media. Usually, documents are duplicated, compiled, or moved across spaces in a shared drive or an enterprise collaboration tool, where the default visibility settings differ. Showing labels such as ”Original sharing scope: Your team — Feb 2024” and allowing one-click carry-over ease coordination failures and reduce interpersonal friction (Lampinen et al., 2011; Petronio, 2002). Narrow-first defaults and clear boundary markers protect against overexposure, particularly when materials persist for long periods, in LMSs or patient portals, where overexposure costs can be much greater.
Privacy labels are in order. Uncertainty detection should transpire and be discarded after gating strictly on the device; maintain only aggregate counters pertaining to evaluating interventions with strict opt-in consent and minimization (Tsai et al., 2011; Leon et al., 2012). Pattern libraries should articulate the rationale where escalations are concerned and should not use deceptive pre-selection or any architecture that nudges the user toward accepting a broader scope of use without their informed consent (Gray et al., 2018; Mathur et al., 2019). Foreground provenance where it is known; otherwise, plainly state the lack of provenance and default conservatively. If the provenance is sensitive (for instance, a prior share to a stigmatized group), its membership should not be revealed in an interface absent explicit user choice; instead, the badge can abstract to a neutral label such as ”Limited audience—May 2024” with an affordance to reveal details.
@@ 3.3. Edge cases
Not all content or communities experience temporal drifting in a similar manner. Some items have unusually high downstream sensitivity (e.g., health-related postings, posted identities with stigma attached, content that was created and shared under distinct norms) such that even a minor inflation of scope can yield very high social costs. For these items, provenance badges should be designed as unequipped with the explicit labels of those items unless a user opts to broaden them, starting with something neutral (i.e., ”Limited audience - May 2024”) and then allowing the user to remove that constraint if they so wish. Defaults for restoration prompts should similarly follow the most narrow viable defaults, with an explicit reverse path toward broadening (Petronio, 2002; Johnson and Goldstein, 2003; Thaler and Sunstein, 2008). Another potential failure mode would happen when a user is predictably incorrect and overconfident; when the user is overconfident, it can overshadow misrecall, simple users should simply nudge backwards along their last recalled norm (like, ”This was originally shared with Friends, keep that as is, or broaden it?”) , and once again avoid silently compliant relationships with users (Koriat, 2012; Moore and Healy, 2008). Finally, designers should take adversarial misuse into account; e.g, if a malicious actor were to strip provenance and re-upload the content causing a default to go public. Defensive heuristics included Narrow-first, for any acceptable post that does not have verifiable provenance within the universe, highlighting user uncertainty towards the poster and supporting single-touch reporting options across suspicious resharing flows (Madejski et al., 2012; Gray et al., 2018; Mathur et al., 2019).
@@ 4. Risk Model and Evaluation for Temporally Aware Privacy UX
Without that ground-truth label of ”correct recall,” the model sells for useful purposes. The teams are able to infer proxies: if there existed previously an audience distinct from the outgoing one (an event of scope expansion), if there is a high amount of hesitation on Old-stuff, or complaints coming downstream in the form of tickets about unintended exposure. If instrumented well and privacy-by-design is baked in, the team can explore several operating points by adjusting an uncertainty threshold and watching aggregate-level guardrails such as completion rate, time-to-post, or override share. It is important that such calibration is done: if predictive scoring is used to rank events by risk, calibration ought to be such that predicted frequencies track true frequencies (Brier, 1950; Niculescu-Mizil and Caruana, 2005; Naeini et al., 2015).
@@
Since experimental access or long-running A/B tests might be infeasible, one could consider an evaluation framework relying solely on the log. First, attach machine-readable provenance wherever possible to the content (for instance, through a lightweight PROV profile (Moreau and Missier, 2013)). Second, add three opt-in counters to measure: (1) how often outgoing scope is strictly broader than attached provenance (used as a conservative proxy for overexposure); (2) override rates (i.e., how often posters choose to broaden or narrow); and (3) completion and time-to-post. Third, simulate counterfactual policies by replaying event logs: identify events that would have triggered a badge or sheet under a candidate policy and apply literature-informed priors for $\tilde{p}_{c},\tilde{p}_{o}$ to estimate $\Delta r$. Fourth, report sensitivity bands instead of point estimates to reflect uncertainty, and prefer policies that dominate across a range of plausible priors.
@@
Scenario Events $p_{c}$ $p_{o}$ $\mathbb{E}[\text{overexposures}]$
Baseline (illustrative) 360 0.57 0.57 $360\times 0.43\times 0.57\approx 88.2$
Prompt top 30% 108 0.72 0.47 $108\times 0.28\times 0.47\approx 14.2$
Unprompted 70% 252 0.57 0.57 $252\times 0.43\times 0.57\approx 61.7$
Total with policy 360 $\approx 75.9$  ( $\downarrow$ 12.3; $\sim$14%)
Table 2. Illustrative cumulative impact under conservative reuse rates. Values are hypothetical.
Instrumenting the patterns seems possible in an ethical manner. Uncertainty proxies (e.g., dwell time over the audience-control after an Old cue) can be calculated locally and discarded. Aggregate counters to support evaluation can maintain their counts in differential-privacy bins or coarser time windows. The product requirement is straightforward: to inform people of what the system knows relevant to the decision (the original audience), so that they may accept, broaden, or narrow with very little ceremony. Governance-wise, it is equally simple: document when prompts appear and for what purpose, and then check to ensure pattern libraries do not quietly broaden scope or hide narrow options (Gray et al., 2018; Mathur et al., 2019).

### v2

# Temporal Drift in Privacy Recall: From Verbatim Loss to Gist-Based Overexposure

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Human-Engaged Computing (ICHEC 2025); November 21–23, 2025; Singapore Management University, Singaporeisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Social aspects of security and privacyccs: Human-centered computing Empirical studies in HCIccs: Security and privacy Usability in security and privacyccs: Human-centered computing HCI design and evaluation methods
@@ 1. Introduction
Figure 1. Temporal drift schematic: recall specificity declines as perceived age increases, raising the chance of scope misremembering.
A simple line chart showing a gentle decline in recall accuracy as perceived age increases.
@@ 1.2. Problem and Stakes
People reshare or cross-post older content all the time. The audience selected at creation might be forgotten months later as context changes. This gap throughtemporal drift in privacy recall opens up unintended disclosure risks, complementing classic context collapse (boyd2011tweet; Vitak2012ContextCollapse). We focus on a ubiquitous visual cue: perceived age-and draw from it design implications for making past audiences recognizable at the moment of reuse.
@@ 1.4. Scope and Assumptions
  • RQ1: How can interfaces convert privacy recall into recognition during reuse with minimal friction?
  • RQ2: How can teams evaluate such patterns without new user studies, using log-only, privacy-preserving methods?
  • RQ3: Why does perceived age reduce accurate recall of past audiences?
  • RQ4: When recall is wrong, why are errors directionally biased toward broader exposure?
1.4.1. Claim scope.
In this paper we treat recall errors tend to bias the other way, biased toward topical over-exposure. This bias is theoretically supported by source-monitoring and gist-over-verbatim poetry aspects, but we are not claiming this is an unbreakable law. In practice, teams using the bias should treat it as a testable prior rather than a fact: when priors are uncertain, the optimal policy is to set bounds conservatively and retain sensitivity bounds.
@@ 2. Related Work
In the literature on social media, the audience reasoning is entwined with the idea of an imagined audience: a poster forms models of the sort of people who might see their content and then behave accordingly-if those mental models are often miscalibrated (boyd2011tweet; Vitak2012ContextCollapse; Marwick2011ImaginedAudience; Litt2016ImaginedAudience; Bernstein2013InvisibleAudience). Due to context collapse, heterogeneous groups of people-family, friends, work colleagues, etc.-are compressed into single broadcast venues, which makes it difficult to maintain types of scopes that must be occasionally nuanced. Even if technological aids are in play, e.g., friends lists, groups, restricted audiences, they demand one to keep up their attention just in time, often at the cost of gradual degradation due to changes in membership or defaults. It is, thus, a social as well as interface problem: the UI lets one choose to emphasize certain dimensions (such as elapsed time) and glaringly omit prior audience, leading to the formation of misinformed choices and decisions (Rader2014Tracking; Eslami2016FolkTheories).
Contextual integrity frames privacy as appropriate information flow—what attributes, transmitted under which principles, to which roles, in which contexts (Nissenbaum2004ContextualIntegrity). Audience provenance—“who could see this then”—is a core parameter of that flow. When resurfacing elides provenance, users must infer appropriateness without a key dimension. That omission is consequential when content migrates across apps with incompatible audience ontologies or when “memories” features highlight time but not the original recipients (Gilbert2023Reuse; Brandtzaeg2018TimeCollapse). Empirically, small changes in defaults and salience can shift disclosure (Wang2014PrivacyNudgesFacebook; Stutzman2010FriendsOnly; Madejski2012PrivacySettings); in reverse, omission of salient parameters can yield miscalibrated behavior.
@@
Imagined & invisible audiences People misestimate who sees posts; mixed groups collapse into one venue (boyd2011tweet; Litt2016ImaginedAudience; Bernstein2013InvisibleAudience). Make the actual prior audience visible at reuse time to correct mental models.
Contextual integrity Privacy = appropriate flows given roles & transmission principles (Nissenbaum2004ContextualIntegrity). Surface “who could see this then” as first-class provenance during reuse.
Memory dynamics Verbatim fades faster than gist; source misattribution (Ebbinghaus1913Memory; Johnson1993SourceMonitoring; Schacter1996Illusions). Convert recall to recognition; show prior scope, not just age.
Choice architecture Small cues & defaults steer disclosure (Wang2014PrivacyNudgesFacebook; Johnson2003DefaultsSaveLives; Thaler2008Nudge). Use narrow-first defaults when provenance is missing; avoid dark patterns.
@@
Memory research makes the temporal dimension sharper. Verbatim details decay rapidly compared to gist (Ebbinghaus1913Memory; Schacter1996Illusions), and people often commit source-monitoring errors—remembering that they shared but misattributing audience or channel (Johnson1993SourceMonitoring; Brainerd2002FuzzyTrace; Roediger1995DRM; Schacter2007RememberingFuture). When a large “6 months ago” label is visible and audience provenance is not, retrieval relies on habits (“I usually post publicly”) or norms (“family photos are for friends”), leading to one-step drift along typical audience orders. Photography and platform automation can further distort memory; for instance, offloading memory to photos or feeds changes what details people recall (Henkel2014Photo; Sparrow2011Google). Overconfidence compounds the problem: people can feel certain while being wrong (Koriat2012SelfConsistencyConfidence; Moore2008Overconfidence).
There need to be three ingredients for temporal drift: the salience of age, invisibility of provenance, and an audience lattice that allows for ”one step more open” errors. Further, some remedies are suggested: Display the relevant parameters at the point of decision (Egelman2010TimingPrivacyIndicators); introduce ethical defaults (Johnson2003DefaultsSaveLives; Thaler2008Nudge); eschew dark patterns (Gray2018DarkPatternsSide; Mathur2019DarkPatterns); and help people form good folk theories by using outcome-based copy, not technical jargon (Rader2014Tracking). We take this into concrete flora of patterns and a risk-aware policy for when to show them.
@@ 3.1. Mechanism & Core Patterns
From the standpoint of temporal drift, it arises as a boundary-case within the recognition-versus-recall framework. If we think about the resurfacing of an old item as a recall event, then it is…” For years, human-computer interaction wisdom has advised favoring recognition over recall because ”recognition means less brain power and fewer chances to go wrong” (Nielsen1994Usability; Sweller1988CognitiveLoad). That seemingly innocent timestamp devoid of provenance violates that principle just when it should have been upheld (Nissenbaum2004ContextualIntegrity; Rader2014Tracking; Eslami2016FolkTheories). The temporary fix is simple: pass the audience parameter so that the user interface offers recognition with possible deferral (Moreau2013PROVDM; Egelman2010TimingPrivacyIndicators). It endorses autonomy as an expression from sharing while making sure there is no unwanted overexposure.
@@
A first pattern is a portable provenance badge: a short, readable label that travels with the media wherever it appears in reuse flows, for example, “Original audience: Friends — May 2024.” The badge answers the precise, situated question posters are likely to have (“who could see this then?”) and can be implemented as metadata read at render time (Nissenbaum2004ContextualIntegrity; Bernstein2013InvisibleAudience; boyd2011tweet; Moreau2013PROVDM). Tapping reveals a small, accessible sheet that preselects “Carry over Friends” and provides one-tap override to broaden or narrow (Johnson2003DefaultsSaveLives; Thaler2008Nudge). The interaction is designed for recognition, not search; it keeps routine reuse fast while making risk-relevant choices reversible and explicit (Nielsen1994Usability).
@@
A second pattern involves the context-restoration prompt, which surfaces only when it is needed most. Many platforms, ranging from WeChat to Instagram, can locally infer indecision; for instance, by detecting on-device hesitation or backtracking through the interface, but without logging raw behavior (Tsai2011PrivacySealsISR; Leon2012OBANotices). When an item is marked as Old, and uncertainty is detected - or if the content category is sensitive - the UI jumps up briefly to confirm its prior audience with a narrow-first alternative (Brandtzaeg2018TimeCollapse). The copy never scolds or talks above the user in jargon but instead describes outcomes (”only your friends will see this”) (Rader2014Tracking; Eslami2016FolkTheories). When user certainty appears to be very high, the UI stays silent unless the requested scope conflicts with provenance, in which case a gentle clarification pops up, with an easy option to preserve the original scope (Koriat2012SelfConsistencyConfidence; Moore2008Overconfidence).
@@
The third pattern comes in for imports without provenance. Cross-app reuse frequently strips or reinterprets audience information (Gilbert2023Reuse). In these situations, it is more ethical to start with the conservative scope while providing an easy way to broaden it than to widen the scope silently. Defaults do possess power; hence, they should be used to defend a user’s rights, not to prey on them (Johnson2003DefaultsSaveLives; Thaler2008Nudge). When audience mappings turn out to be lossy or ambiguous, the interface should make the ambiguity explicit and let the poster decide (Gray2018DarkPatternsSide; Mathur2019DarkPatterns).
@@ 3.2. Accessible Deployment, Portability, & Governance
Those patterns have to be accessible and international in implementation. Screen-reader labels should expose the badge text (e.g., ‘Original audience: Friends‘) and the purpose for controls (e.g., ‘carry over original audience‘). Labels, icons, and focus order must allow keyboard navigation; color contrast must consider common impairments; date formats must be localized; and right-to-left layouts need validation. Copies are outcome-oriented, not mechanism-oriented. They say what will happen (‘only your friends will see this‘) and not how the system judges it. Because users build folk theories about feeds and privacy controls (Eslami2016FolkTheories; Rader2014Tracking), correct and easy-to-understand explanations reduce the confusion and increase the users’ trust.
The patterns are also portable beyond social media. Usually, documents are duplicated, compiled, or moved across spaces in a shared drive or an enterprise collaboration tool, where the default visibility settings differ. Showing labels such as ”Original sharing scope: Your team — Feb 2024” and allowing one-click carry-over ease coordination failures and reduce interpersonal friction (Lampinen2011Together; Petronio2002BoundariesOfPrivacy). Narrow-first defaults and clear boundary markers protect against overexposure, particularly when materials persist for long periods, in LMSs or patient portals, where overexposure costs can be much greater.
Privacy labels are in order. Uncertainty detection should transpire and be discarded after gating strictly on the device; maintain only aggregate counters pertaining to evaluating interventions with strict opt-in consent and minimization (Tsai2011PrivacySealsISR; Leon2012OBANotices). Pattern libraries should articulate the rationale where escalations are concerned and should not use deceptive pre-selection or any architecture that nudges the user toward accepting a broader scope of use without their informed consent (Gray2018DarkPatternsSide; Mathur2019DarkPatterns). Foreground provenance where it is known; otherwise, plainly state the lack of provenance and default conservatively. If the provenance is sensitive (for instance, a prior share to a stigmatized group), its membership should not be revealed in an interface absent explicit user choice; instead, the badge can abstract to a neutral label such as ”Limited audience—May 2024” with an affordance to reveal details.
@@ 3.3. Edge cases
Not all content or communities experience temporal drifting in a similar manner. Some items have unusually high downstream sensitivity (e.g., health-related postings, posted identities with stigma attached, content that was created and shared under distinct norms) such that even a minor inflation of scope can yield very high social costs. For these items, provenance badges should be designed as unequipped with the explicit labels of those items unless a user opts to broaden them, starting with something neutral (i.e., ”Limited audience - May 2024”) and then allowing the user to remove that constraint if they so wish. Defaults for restoration prompts should similarly follow the most narrow viable defaults, with an explicit reverse path toward broadening (Petronio2002BoundariesOfPrivacy; Johnson2003DefaultsSaveLives; Thaler2008Nudge). Another potential failure mode would happen when a user is predictably incorrect and overconfident; when the user is overconfident, it can overshadow misrecall, simple users should simply nudge backwards along their last recalled norm (like, ”This was originally shared with Friends, keep that as is, or broaden it?”) , and once again avoid silently compliant relationships with users (Koriat2012SelfConsistencyConfidence; Moore2008Overconfidence). Finally, designers should take adversarial misuse into account; e.g, if a malicious actor were to strip provenance and re-upload the content causing a default to go public. Defensive heuristics included Narrow-first, for any acceptable post that does not have verifiable provenance within the universe, highlighting user uncertainty towards the poster and supporting single-touch reporting options across suspicious resharing flows (Madejski2012PrivacySettings; Gray2018DarkPatternsSide; Mathur2019DarkPatterns).
@@ 4. Risk Model and Evaluation for Temporally Aware Privacy UX
We formalize gating as a cost-sensitive decision over features $\mathbf{x}$ that yield estimates of $p_{c}$ and $p_{o}$, combined into a risk score $(1-p_{c})p_{o}$, and compare expected benefit $c_{o}\Delta r$ to friction $c_{f}$; see Fig. 2.
[…]
Figure 2. Cost-sensitive gating: estimate $p_{c}$ and $p_{o}$ to form $(1-p_{c})p_{o}$, then intervene when the risk-weighted benefit $c_{o}\Delta r$ exceeds friction $c_{f}$.
4.0.1. Decision-analytic model.
@@
4.0.2. Estimating $p_{c}(\mathbf{x})$ and $p_{o}(\mathbf{x})$.
Without that ground-truth label of ”correct recall,” the model sells for useful purposes. The teams are able to infer proxies: if there existed previously an audience distinct from the outgoing one (an event of scope expansion), if there is a high amount of hesitation on Old-stuff, or complaints coming downstream in the form of tickets about unintended exposure. If instrumented well and privacy-by-design is baked in, the team can explore several operating points by adjusting an uncertainty threshold and watching aggregate-level guardrails such as completion rate, time-to-post, or override share. It is important that such calibration is done: if predictive scoring is used to rank events by risk, calibration ought to be such that predicted frequencies track true frequencies (Brier1950Verification; NiculescuMizil2005Predicting; Naeini2015BBQ).
@@
Since experimental access or long-running A/B tests might be infeasible, one could consider an evaluation framework relying solely on the log. First, attach machine-readable provenance wherever possible to the content (for instance, through a lightweight PROV profile (Moreau2013PROVDM)). Second, add three opt-in counters to measure: (1) how often outgoing scope is strictly broader than attached provenance (used as a conservative proxy for overexposure); (2) override rates (i.e., how often posters choose to broaden or narrow); and (3) completion and time-to-post. Third, simulate counterfactual policies by replaying event logs: identify events that would have triggered a badge or sheet under a candidate policy and apply literature-informed priors for $\tilde{p}_{c},\tilde{p}_{o}$ to estimate $\Delta r$. Fourth, report sensitivity bands instead of point estimates to reflect uncertainty, and prefer policies that dominate across a range of plausible priors.
@@
Scenario Events $\mathbf{p_{c}}$ $\mathbf{p_{o}}$ $\mathbb{E}[\text{overexposures}]$
Baseline (illustrative) 360 0.57 0.57 $360\times 0.43\times 0.57=88.24$
Prompt top 30% 108 0.72 0.47 $108\times 0.28\times 0.47=14.21$
Unprompted 70% 252 0.57 0.57 $252\times 0.43\times 0.57=61.78$
Total with policy 360 0.62 0.54 $14.21+61.78=\mathbf{76.00}$
Table 2. Illustrative cumulative impact under a targeted prompting policy.
Instrumenting the patterns seems possible in an ethical manner. Uncertainty proxies (e.g., dwell time over the audience-control after an Old cue) can be calculated locally and discarded. Aggregate counters to support evaluation can maintain their counts in differential-privacy bins or coarser time windows. The product requirement is straightforward: to inform people of what the system knows relevant to the decision (the original audience), so that they may accept, broaden, or narrow with very little ceremony. Governance-wise, it is equally simple: document when prompts appear and for what purpose, and then check to ensure pattern libraries do not quietly broaden scope or hide narrow options (Gray2018DarkPatternsSide; Mathur2019DarkPatterns).
4.0.3. Operating under uncertainty.
If directionality is uncertain, we can optimize on upper bounds rather than point estimates. Let $p_{c},p_{o}$ lie in intervals $l_{c},u_{c}],[l_{o},u_{o}]$ derived from literature-informed Beta priors or log-only proxies. Then a safe, cost-sensitive trigger compares the worst-case risk
$\mathbb{E}[R]_{\max}=(1-l_{c})\,u_{o}$
against friction $c_{f}$ via $c_{o}\cdot\Delta r_{\min}>c_{f}$, where $\Delta r_{\min}$ uses conservative bounds on intervention lift $(\tilde{p}_{c}-p_{c})$ and $(p_{o}-\tilde{p}_{o})$. This yields a policy that reduces overexposure whenever it is plausibly present, without assuming its magnitude.